HUBBLE SPACE TELESCOPE DAILY REPORT #818 PERIOD COVERED: 0800 EST 02/05/93 - 0800 EST 02/08/93 Daily Status Report as of 039/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 1160 (Absorption Cl. Phys. in Seyfert Galactic Nuclei) The High resolution Spectrograph was used to make fifteen observations of the nucleus of Seyfert galaxy NGC-3783. All fifteen observations were reported to look okay. 1.2 Completed GHRS/2 4012 (Dark Count Statistics for GHRS Detectors) The High Resolution Spectrograph was used to make one dark observations for a special investigation of the number 2 detector. The single observation was received and displayed, and was reported to look okay. 1.3 Completed Two Sets of GHRS/2 3614 (Boron as a Probe of Stellar Structure) The High Resolution Spectrograph was used to make Boron line studies of the star HD26462 to study the star's interior. The guide star acquisition failed. MOSES uplinked an Attitude Reference Update and the reacquisition was successful. An HSTAR was written. An HSTAR was also written for a carousel reset error that was received prior to the first observation. The second set of science observations were received and were reported to look "pretty good", but it is not sure if the target was well centered in the aperture. 1.4 Completed WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to take two internal flat observations using filter F555W looking for daisies and measles (i.e., new form of contamination). Both to the observations were reported to be okay. 1.5 Completed PC 3230 (PC PSF and Focus Monitoring 2) The Planetary Camera was used to make one image of the point spread function and to test the system focus by making an observation of quasar SAO55685. The image was reported to have fallen on the Baum Spot. 1.6 Completed HSP/UV1 4247 (UV Observations of Geminga) The High Speed Photometer was used to make an ultraviolet study the optical counterpart of the gamma ray source Geminga (aka, GSC55685-0). The object was reported to be well centered, but exceedingly faint, as was expected. 1.7 Completed Two Sets of WFC/PC 3172 (WF/PC SV Observations: Internal Monitor Proposal) Both the Wide Field Camera and the Planetary Camera were used to take a series of dark, internal flat, and KSPOT images to be used in the calibration of the camera. All observations were reported to be normal calibration exposures. 1.8 Completed PC 4347 (WF/PC Calibration: Flood Retention Test) The Planetary camera was used to that three Earth flat observations using filter F439W to determine how much of the last ultraviolet flood retention test still remains. All three observations were received and displayed. The first image was underexposed, and the last two were overexposed. 1.9 Completed PC 4712 (WF/PC Calibration: Measles Monitor) The Planetary Camera was used to obtain Earth flats using filter F517N, both with and without filter F122M. This is to monitor the appearance and number of measles (i.e., a form of contamination). Three image were received and displayed. The first image was underexposed, and the other two were reported to be okay. 1.10 Completed GHRS/2 3626 (Ephemeral Deter. of Wind Velocity and Density) The High Resolution Spectrograph was used to make thirteen spectral observations of the eclipsing binary star Zeta Aurigae (aka, HD32068). All of the calibration images and science spectral observations were reported to be okay. 1.11 Completed GHRS/2 3737 (the Onset of Chromospheric Activity) The High Resolution Spectrograph was used to make four spectral observations of the area around the carbon II resonance lines at 1335 A0 on A class star HD165040. All four observations were executed without problems. The science data was reported to be very noisy, and possibly underexposed. 1.12 Completed FOS/BL 3268 (UV Spectra of QSOs with Z > 3.1 Observations) The Faint Object Spectrograph was used to make three ultraviolet spectral observations of high red shifted quasar 1334-005. All three observations were executed without problems, and the science data was reported to be okay. 1.13 Completed WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera was used to take two high latitude images which were taken in parallel with the prime science observations. The images were reported to look okay. 1.14 Completed GHRS/2 3934 (Physical Conditions and Velocity Structures) The High Resolution Spectrograph was used to make spectral observations of hot red giant binary star CI Cygni. The base line guide star acquisition failed coarse track. A HSTAR was written. The subsequent two reacquisitions also failed. The subsequent science observations were reported as being lost. Later a third reacquisition was successful. There were no observations made after this, since the HRS schedule to stop observing was less than 15 minutes after the reacquisition. 1.15 Completed FOS/RD 3840 ( Abundances and Evolution of Carbon Nitrogen and Oxygen) The Faint Object Spectrograph was used to make ultraviolet spectra of the extra galactic hydrogen II region T1214-277 looking for the abundances of carbon, nitrogen, and oxygen. The observations were executed as scheduled, and no problems were reported. 1.16 Completed WFC 3917 (HST Medium Deep Survey Part 1) The Wide Field Camera was used to take a series of high latitude survey images which were taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 1.17 Completed FGS 2941 (High Speed Astrometry Search for Planet Companions) The FGSs were used to make astrometric observations of Barnard's star looking for the presence of a Jupiter sized planet. The observations were monitored in the OSS, and no problems were reported. 1.18 Completed FOC 3790 (Polarization Imaging of Radio Galaxies) The Faint Object Camera (f/96)was used to make three ultraviolet polarization image of radio galaxy 3C321. All three images were received and displayed in the OSS. They were reported to look okay. 1.18 Completed FOC/96 3801 (QSOs Suitable for Subsequent Observations) The Faint Object Camera (f/96) was used to make one image of quasar 1640+4628 to measure its ultraviolet flux level to see if it has sufficient flux to be observed in future observations using the GHRS and FOS. 1.20 Completed PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to take one image of the protoplanetary system HE3-1475. The observation was taken during LOS, and still remains on the science tape recorder. It will be analyzed later. 1.21 Initiated FOC/96 3487 (Cooling Flows in Distant Clusters) The Faint Object Camera (f/96) is currently being used to make observations of the star formation region in galaxy NGC-5920. The observations are currently on-going, and will be analyzed later. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 15 Successful : 13 (037/0245Z AND 038/0046Z) Scheduled Reacquisitions : 19 Successful : 17 038/0219Z AND 038/0355Z) Losses of Lock : 0 2.2 FHST Updates: Scheduled : 37 Successful : 36 (037/2315Z) 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 3487 (Cooling Flows in Distant Clusters) 3.2 Complete GHRS/2 3614 (Boron as Probe of Stellar Structure and Mass Loss) 3.3 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.4 Uplink Updated Ephemeris Tables, Per 2871 3.5 Complete FOC/96 3801 (Search for QSOs Suitable for Subsequent Observations) 3.6 Complete GHRS/2 4223 (Absorption cross Sections of Nearby Galaxies) 3.7 Initiate GHRS/2 2537 (Physical Conditions of Interstellar Gas) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-